Use the given information to find the values of sin α/2 , cos α/2 , tan α/2
sin α = 8/17 ; π/2 < α < π


a. sin α/2 = ___
b. cos α/2 = ___
c. tan α/2 = ___

Answer by stanbon(75887)

Use the given information to find the values of sin α/2 , cos α/2 , tan α/2
sin α = 8/17 ; π/2 < α < π
Find cos(a)::
Since sin(a) = 8/17, y = 8, r = 17
So x = sqrt(17^2-8^2) = 25
And cos(a) = x/r = 25/17
--------------
a. sin α/2 = sqrt[(1-cos(a))/2] = sqrt[(1-25/17)/2]
b. cos α/2 = sqrt[(1+cos(a)/2] = sqrt[(1+(25/17))/2]
c. tan α/2 = sqrt[(1-(25/17)/1+(25/17)]
--------------
Cheers,
Stan H.
--------------------